OpenStudy (kendricklamar2014):

A glass cube has edges that are 3/4 of the length of the edges of a cube. The larger cube has a surface area of 6,000 cm^2. What is the surface area of the smaller cube? The Ratio of the surface area is _______ How can you use the ratio of the surface areas in a proportion to find the surface area of the smaller cube?

OpenStudy (anonymous):

get the surface area of smaller cube using the formula : surface area =6*a^2 where a is the edge length. then multiply it with (4/3) to get edge of larger one then calculate the s.area of larger one. Divide it with surface area of smaller one to get answer

OpenStudy (mathmale):

Kendrick: What have you done so far? Have you tried drawing the two cubes and labeling their edges? Have you asked yourself what the formula for the total surface area of a cube is?

OpenStudy (mathmale):

If you know that the surface area of the larger cube is 6000 square cm, then what is the length of each side of the larger cube? Of the smaller cube?

OpenStudy (kendricklamar2014):

is it 6,000/8?

OpenStudy (mathmale):

Kendrick, a cube has only 6 sides, not 8. You'd possibly find this problem to be easier if you'd please write out the formula for the surface area of a cube in terms of the length of any one side of that cube.

OpenStudy (kendricklamar2014):

6,000/6= 1000

OpenStudy (mathmale):

That's better. What does that 1000 represent? Remember, you need to find the length of each side of the cube, and that your 1000 is only an intermediate result (not the answer).

OpenStudy (mathmale):

Kendrick: Please find and type here the formula for the total surface area of a cube. Thanks.

OpenStudy (anonymous):

@KendrickLamar2014 Please do some online research for formula of the geometrical figures, note it down and learn by heart.

OpenStudy (kendricklamar2014):

A=6a^2

OpenStudy (mathmale):

Kendrick, you can do this problem! If you're familiar with Internet searches, Google "area of a cube." Please, keep going. Yes, the surface area of a cube is A = 6 a^2, where a is the length of any edge of the cube in question. Nice work!

OpenStudy (mathmale):

Repeat: The area of the larger cube is 6000 square centimeters. What is the length of one side of this cube? Hint: Use that formula you found earlier.

OpenStudy (kendricklamar2014):

1,000?

OpenStudy (mathmale):

Total surface area = 6000 square cm. Surface area of ONE side of the cube = (6000 square cm)/6 = 1000 square cm. Are of a square of side length " a " is A = a^2. If A = 1000 square cm, what is the length of one of the sides of this square area?

OpenStudy (kendricklamar2014):

1,000?

OpenStudy (kendricklamar2014):

@mathmale is it 1,00

OpenStudy (kendricklamar2014):

i mean 1,000

OpenStudy (kendricklamar2014):

@johnweldon1993 can you help me please